# Video Understanding with Large Language Models: A Survey

### v4

@@ V-B Interactive and User-Centric Technologies
  • State-Aware Human-Computer Interaction and Robot Planning: In the field of human-computer interaction (HCI), Vid-LLMs analyze user videos to discern context and provide customized assistance, as highlighted in Bi et al. [322]. Interaction forms also involve video content understanding like captioning videos [323, 324, 149]. Concurrently, in autonomous robot navigation, the SayPlan method [325] integrates LLMs with 3D scene graphs to enable robots to interpret and navigate complex spaces in large buildings.
V-C Healthcare and Security Applications
@@
  • Healthcare Innovations: In the healthcare sector, Vid-LLMs play a crucial role in processing and interpreting medical literature, assisting in diagnostic and educational processes [326, 327, 328, 329], and providing decision support for healthcare professionals. They are utilized in patient interaction tools, such as chatbots for symptom assessment and addressing health-related queries, thereby improving patient care and access to information [330].
@@
  • Security, Surveillance, and Cybersecurity: Vid-LLMs are crucial in security and protection, analyzing communications for potential threats [331, 332] and detecting anomalous patterns in data [333, 334]. In surveillance video analysis, they identify suspicious behaviors, aiding law enforcement [335]. Their role in cybersecurity includes identifying phishing attempts and contributing to forensic analysis by summarizing case-related texts [336].
@@
  • Advancements in Autonomous Vehicles: In autonomous vehicles, Vid-LLMs process natural language inputs for interaction [337], assist in understanding road signs and instructions [241, 338], and improve user interfaces for vehicle control systems [337], enhancing safety and user experience.
VI Future Directions and Conclusion
In this survey, we have reviewed the latest advancements in video understanding methods with large language models (Vid-LLMs) and introduced the fundamental principles, significant discoveries, and techniques for understanding and applying Vid-LLMs effectively. We begin with the developmental history of video understanding, from traditional non-deep learning methods to neural network-based approaches, then self-supervised pretraining for videos, and now the current video understanding solutions based on LLMs.
VI-A Limitations and Future Work
Although current methods enhance the effectiveness of various downstream tasks by pretraining and fine-tuning with a large amount of video data, and the introduction of LLMs allows for a better understanding of various information in videos, there remain many challenges unresolved when facing diverse video understanding tasks in the real world.
  • Fine-grained Video Understanding: Fine-grained video understanding remains a challenge. Whether it’s in the field of temporal understanding of videos or visual grounding, the lack of datasets and insufficient research has made it difficult to achieve finer granularity in video understanding tasks. Also, processing and analyzing video data requires significant computational resources. Fine-grained understanding typically means analyzing every video frame, significantly increasing the computational load. Furthermore, videos contain not just spatial information but also temporal information. Understanding how objects change and interact over time, especially at a fine-grained level, is much more complex than expected. A step further, understanding the deeper semantics of video content, such as emotions, metaphors, or the dynamics of complex scenes, is more difficult than merely recognizing objects or actions. The combination of LLMs and videos has brought a ray of hope for fine-grained video understanding. LLMs enable text to align with videos at various levels, partially addressing the issue of fine-grained video understanding.
@@ VI-A Limitations and Future Work
  • Long-form Video Understanding: Since long videos contain vast amounts of frames, the extended duration of long videos adds complexity to the analysis, especially in understanding events and behaviors over time. Thus, identifying key events and maintaining attention in long videos is difficult. Effective mechanisms are needed to detect and highlight important parts, particularly in content-rich or complex plot videos.
@@
  • Multimodal Video Understanding: Multimodal video understanding requires the integration of different types of data, such as visual, audio, and text, for a better understanding of videos. Aligning these data, especially in terms of spatial and temporal synchronization, is particularly critical. This area lacks relevant research and suffers from a scarcity of datasets. Moreover, constructing such datasets faces challenges as ensuring high quality and consistency in data annotation is often difficult. In terms of methodology, extracting and utilizing effective features across different modalities is key to achieving precise video understanding, but this process is filled with challenges.
  • Human Interaction in Video Understanding: The results of video understanding ultimately serve humans, so how to better convey human needs and understand model results is also a very important issue. The emergence of LLMs has enabled video understanding models and humans to convey information more effectively through text. However, LLMs have not completely solved the interaction problem. For example, using text to guide the model’s understanding of a video cannot handle extremely fine-grained video understanding; also, when the model outputs text, it cannot precisely describe the complex content in the video. In addition, due to the capabilities of the video encoder and LLMs, some high-level information, such as character emotions and attitudes, cannot be well represented. Therefore, how to use other prompts, such as points, scribbles, etc., to optimize the interaction between humans and video understanding models is worth studying. In addition, how to improve the video encoder’s ability to preserve details is also a major issue.
  • Hallucination in Multimodal LLMs: “Hallucination” refers to the phenomenon where the model generates responses that are significantly disconnected from the relevant source material or input. This can lead to the creation of highly erroneous or unrealistic descriptions that do not align with the provided videos. In video understanding with LLMs, the main reasons for this situation are as follows: 1) Insufficient extraction of visual features. 2) The influence of video contextual content. 3) The domain gap between the visual feature domain and the Language domain. 4) Hallucination inherent in LLMs. Therefore, to address the impact of hallucination, solutions can be found in improving the effectiveness of video encoders, enhancing the understanding of long-form spatio-temporal contexts, and the collaboration between visual latent space and linguistic latent space.
VI-B Conclusion
@@
In this survey, we have thoroughly examined the advancements and applications of large language models (LLMs) in the domain of video understanding. The integration of LLMs into video analysis has marked a significant leap forward, enabling more sophisticated and versatile processing capabilities that surpass traditional methodologies. We categorized the current approaches into three main types: Video Analyzer $\times$ LLM, Video Embedder $\times$ LLM, and (Analyzer + Embedder) $\times$ LLM, with further sub-classifications based on the functional roles of LLMs such as Summarizer, Manager, Text Decoder, Regressor, and Hidden Layer. The survey underscores the emergent capabilities of Vid-LLMs in tasks requiring multi-granularity reasoning, from abstract to spatiotemporal analysis, demonstrating their potential in diverse applications such as video summarization, captioning, question answering, and more. Despite the significant progress, we also identified several limitations and challenges, including the need for more robust evaluation metrics, handling long-form videos, and improving the alignment between visual and textual modalities. Future research directions are poised to address these challenges by exploring more efficient training strategies, enhancing the scalability of Vid-LLMs, and developing innovative architectures that further integrate multimodal data. Additionally, expanding the scope of datasets and benchmarks will be crucial in driving the next wave of advancements in video understanding with LLMs. Overall, the continued evolution of Vid-LLMs holds promise for transforming video analysis into a more automated, accurate, and contextually aware process, paving the way for numerous practical applications in various fields.

### v5

@@ V-A2 Interactive and User-Centric Systems
  • State-Aware Human-Computer Interaction and Robot Planning: In the field of human-computer interaction, Vid-LLMs analyze user videos to discern context and provide customized assistance, as highlighted in Bi et al. [340]. Interaction forms also involve video content understanding like captioning videos [341, 342, 155]. Concurrently, in autonomous robot navigation, the SayPlan method [343] integrates LLMs with 3D scene graphs to enable robots to interpret and navigate complex spaces in large buildings.
V-A3 Healthcare and Security Applications
@@
  • Healthcare Innovations: In the healthcare sector, Vid-LLMs play a crucial role in processing and interpreting medical literature, assisting in diagnostic and educational processes [344, 345, 346, 347], and providing decision support for healthcare professionals. They are used in patient interaction tools, such as chatbots for symptom assessment and addressing health-related queries, thus improving patient care and access to information [348].
@@
  • Security, Surveillance, and Cybersecurity: Vid-LLMs are crucial in security and protection, analyzing communications for potential threats [349, 350] and detecting anomalous patterns in data [351, 352]. In surveillance video analysis, they identify suspicious behaviors, helping law enforcement [353]. Their role in cybersecurity includes identifying phishing attempts and contributing to forensic analysis by summarizing case-related texts [354]. They may also improve video crowd counting [355] for security applications.
@@
  • Autonomous Vehicles: In autonomous vehicles, Vid-LLMs can process language inputs for interaction [356], assist in understanding road signs and instructions [247, 357], and improve user interfaces for vehicle control systems [356], enhancing safety and user experience.
V-A4 Other Applications
Vid-LLMs offer valuable applications beyond those previously discussed. In video generation research [358, 359, 360], Vid-LLMs can evaluate model performance, refine text prompts, and provide reasoning capabilities that better reflect human intentions. Additionally, Vid-LLMs show promise in resource-constrained environments through edge computing applications [361, 362, 363] and can enhance privacy-preserving distributed systems through federated learning frameworks [364, 365, 366].
V-B Future Directions
Despite enhancing multiple downstream tasks, Vid-LLMs face several challenges in real-world video understanding:
V-B1 More Fine-grained Video Understanding
Fine-grained understanding remains challenging due to limited datasets, insufficient research, and high computational demands. The frame-by-frame analysis increases computational load while capturing spatiotemporal information. Understanding deeper semantics (emotions, scene dynamics) is harder, though text-video alignment through LLMs offers promise [292].
V-B2 Long-form Video Understanding
@@
Long videos’ extended duration complicates the analysis, especially in understanding events over time. Thus, identifying key events and maintaining attention in long videos is difficult [196, 213, 186]. Effective mechanisms are needed to detect and highlight important parts, particularly in content-rich or complex plot videos.
V-B3 Multimodal Video Understanding
@@
Multimodal video understanding requires integrating different types of data, such as visual, audio, and text, for a better understanding of videos [120, 367]. Aligning these data, especially in terms of spatial and temporal synchronization, is particularly critical. This area lacks relevant research and suffers from a scarcity of datasets. The field lacks research and datasets, with challenges in ensuring high-quality data annotation.
V-B4 Hallucination in Video LLMs
Hallucination occurs when models generate responses disconnected from source material [368], caused by insufficient feature extraction, influence of video context, domain gap between vision and language, and inherent LLM hallucinations. Solutions include post-training strategies [210], enhanced spatiotemporal context understanding, and visual-linguistic latent collaboration.
V-B5 Industrial Deployment and Scalability
Effective deployment strategies [213, 369, 370, 212, 371, 362] include model compression, token merging, domain-specific fine-tuning, modular architectures, efficient caching, and standardized integration frameworks, balancing efficiency with performance for industrial systems.
V-C Ethical Implications
The ethical implications of Vid-LLMs center on privacy, data security, and potential misuse. These models perform tasks like video engagement analysis, transcription, summarization, and captioning, requiring access to sensitive content. This raises privacy risks, as video data may contain private or confidential information that could be exposed without proper consent. Also, Vid-LLMs can be misused for surveillance or generating misleading content. Bias is another concern, especially if training data lacks diversity. Addressing these issues requires robust data governance, consent mechanisms, and ethical deployment to prioritize privacy and fairness.
VI Conclusion
@@
This survey has examined the integration of LLMs in video understanding, which has enabled more sophisticated and versatile processing capabilities beyond traditional methods. We categorized current approaches into three main types: Video Analyzer $\times$ LLM, Video Embedder $\times$ LLM, and (Analyzer + Embedder) $\times$ LLM, with sub-classifications based on LLM functional roles: Summarizer, Manager, Text Decoder, Regressor, and Hidden Layer. Vid-LLMs demonstrate capabilities in multi-granularity reasoning from abstract to spatiotemporal analysis, showing potential across video summarization, captioning, question answering, and other applications. Despite progress, limitations remain in evaluation metrics, long-form video handling, and visual-textual modality alignment. Future research will address these challenges through more efficient training strategies, improved Vid-LLM scalability, innovative architectures for multimodal integration, enhanced long-form video understanding, and methods to mitigate hallucinations. Expanding datasets and benchmarks will be critical for advancing video understanding with LLMs.